A machine set to fill soup cans with a mean of 20 ounces and a standard deviation of .11 oz. A random sample of 15 cans has a mean of 20.04 oz. Should the machine be reset?
1
Expert's answer
2020-09-28T20:36:32-0400

solution


From the population, "\mu =20" and "\sigma=0.11"

The sample mean, "\bar x =20.04"



"z=\frac{\bar x - \mu}{\sigma}"

"z=\frac{20.04-20}{0.11}=0.3667"

Since the machine could underfill or overfill the cans, we use a 2 tail distribution. At a 95% level of confidence, "z=1.96" . Therefore we would reject the hypothesis that the two means are equal if the calculated "z\ value > 1.96"


answer:

At "z=0.3667," the samples filled are within the expected limits hence the machine should not be reset